The Nature About Truth and Beauty

Pekka Sinervo University of Toronto

Overview

• What is Inside a Proton?

• Forces, Leptons and Quarks

• Discovery of Beauty

• Discovery of Truth

• Whither Next?

University College 15 March 1999

Inside a Proton – Some High School Physics

Start by reviewing today’s high school science

• Matter is made up of atoms

• There are about 109 different species!

• Atom itself has structure:

– consists of small hard “core” (nucleus) ∗ nucleus made of protons and neutrons

– large cloud of electrons

• We can describe all atoms using 3 particle types 1. electrons (e^{−})

2. protons (p)

3. neutrons (n)

But there are hundreds of particles like p and n’s

• Leptons

• Hadrons

Some History About Structure

Lord Rutherford performed seminal scattering studies

• Looked at α scattering off gold (1911)

• Consistent with Coulomb scattering off point particle

⇒ Uncovered the nucleus

Hofstadter and colleagues extended this work

• Studied 1 GeV e^{−} scattering off nuclei (1953)

• Found that nucleus had extended charge distribution

⇒ shape of the nucleus

Taylor and colleagues probed more finely

• Used ∼ 18 GeV e^{−} scattering off proton (1968)

• Found that nucleons were extended objects ⇒ discovered quarks!

About Rutherford Scattering

Rutherford developed classical theory of scattering

• Charged particle Z_{1}e scattering off nucleus Ze

• Define q⃗ = p⃗ − p′, so that q = 2p sin \frac{θ}{2}

• Differential cross section is

(\frac{dσ}{dΩ} )_{R} = \frac{4m^{2} (Z_{1}Ze^{2})^{2}}{q^{4}}

• Correct classical and QM result ⇒ Rutherford Scattering

Discovered what we now call the nucleus

• To Rutherford, it looked like a point

• We now know it is ∼ 10^{−12} m in size

What Happens for Extended Objects?

When scattering particle is extended in space

• Introduce “form factor” F (q^{2})

F (q^{2}) = ∫ ρ(R⃗)e^{iq⃗·R⃗}d^{3} R⃗

• Fourier transform of “charge” density ρ

⇒ (\frac{dσ}{dΩ} ) = (\frac{dσ}{dΩ} )_{R} |F (q^{2})|^{2}

Proton Structure

Deep inelastic e^{−}p scattering reveals proton structure

• study inelastic part of cross section

e^{−}p → e^{−}X_{hadrons}

– suppress “coherent” and resonant nuclear scattering – Begin to see scattering off of constituents

– Hitting partons with smaller x (x is momentum fraction)

• epitomised by “structure functions” F_{i}(x, q^{2})

(\frac{d^{2}σ}{dq^{2}dx}) ∼ \frac{4πα^{2}}{q^{4}} [(1 − y)\frac{F_{2}(x, q^{2})}{x} + ...]

Page 6 of ?? pages.

Deep Inelastic Scattering Results

SLAC DIS studies revealed clear internal structure

• Evident as hadronic energy W increases (decreases x)

Detailed studies of structure functions F_{1} and F_{2} show

• Charge 1/3 and 2/3 particles

• Spin 1/2 particles

• Presence of anti-particles (the “sea”)

Page 7 of ?? pages.

Bricks and Mortar in 1975

The “Standard Model” of Elementary Particle Physics

• Four “generations” of quarks and leptons

( \begin{matrix}u \\ d \end{matrix} ) ( \begin{matrix}c \\ s \end{matrix} ) ( \begin{matrix}Q = +2/3 \\ Q = −1/3 \end{matrix} )
( \begin{matrix}e \\ ν_{e} \end{matrix} ) ( \begin{matrix}μ \\ ν_{μ} \end{matrix} ) ( \begin{matrix}Q = −1 \\ Q = 0 \end{matrix} )

Come in “pairs” because of the way forces acted on them • Forces carried by bosons

Electromagnetic force  γ
Weak force  W^{±}, Z^{∘} bosons
Strong force  eight gluons
Gravity  gravitons

• Quark and lepton mass due to “Higgs” particle

How this Picture Developed

Leptons appeared fundamental

• electron is the most ubiquitous

• other leptons can be produced and studied

• neutrinos are particularly “ephemeral”

Colour force predicts that quarks only bind when

• three quarks (or three anti-quarks)

• a quark and an anti-quark

For example, a proton is made up of 3 quarks

• 2 up quarks

• 1 down quark

Four flavours ⇒ get all known “elementary particles”

• This model works very well

But Some Exciting Problems

In 1977, team led by Leon Lederman discovered a new particle

• Heaviest particle ever seen (about 10 proton masses)

• Discovered in remnants of very high energy p − p collisions • Looked like a neutral meson made up of a particle and anti-

particle

– Had to be a new, unexpected type of quark

– Called b quark (for either “beauty” or “bottom”)

– Weighed about 5 proton masses

– Production rate implied Q = −1/3 e

What did this mean?

• New colliders at Cornell and Hamburg started studying its properties

• Was heavy, lived for a fairly long time (about 10^{−12} s)

• The way electroweak force acted on it ⇒ Had to have a partner ⇒ t quark

Status of Things in 1985

Some pieces of this puzzle were not there

1. Expected “Unification” of forces • don’t have a clue as to how to understand gravity • only have hints that forces are unified

2. Don’t understand mass at all • “Higgs” boson has not been seen.....

3. Missing the last quark

But the rest of it worked extremely well!

• Many tests have been made over last 20 years

Challenge is to get data to address these questions

• ⇒ Focus on the incomplete pieces

Why MUST the Top Quark Exist?

Other experiments showed that if top doesn’t exist, the theory just dies

• Masses of the W^{±} and Z^{∘} bosons cannot be explained

– Measurements of the Z boson require

M_{top} = 166_{−19−22}^{+17+19} GeV/c^{2}

• Behaviour of the bottom quark implies that it must have a partner

– Weak force predicts asymmetry in e^{+} e^{−} → bb̄

• B meson “mixing” could not occur

Searching for Top

Big Bang was last time nature created top quarks

• Can now pair-produce a heavy quark with enough energy • Require high energy matter-antimatter collisions

• Need tons of data (they are rare events)

pp̄ Collider at Fermilab was the place to search for it

• High luminosities (L ∼ 25 × 10^{30} cm^{−2}s^{−1} )

– compare to radiation in this room L ∼ 2×10^{−2} cm^{−2}s^{−1} )

• Highest energies (√s = 1800 GeV)

– Enough energy to form fairly complex molecule

• Collected 3 million collisions

Why are Heavy Quarks Hard to Find?

Hard to produce

• Must sift through lots of matter-antimatter collisions

• Happens about 1 in every billion collisions?

They Decay!

• A top quark decays via the weak force into a bottom quark • The bottom quark then decays into lighter quarks

Ya can’t see free quarks

• Quarks are “not allowed” to appear unescorted in public • Have to “dress” themselves up around other quarks

• Result in “jets” of particles

Top Quark Signatures

Top quarks only decay via the weak force

t → W^{+} b
\underrightarrow{|}ud̄ or cs̄

\underrightarrow{|}e^{+} ν_{e} or μ^{+} ν_{μ} or τ^{+} ν_{τ}

This results in a number of very unusual “signatures”

• Can use these characteristics to sift the collisions

The Race for the Top

By 1987, competition on to search for the top quark

• Joined an “intimate” collaboration of about 300 physicist • Constructed the CDF detector

– Large 5000 ton detector

– Think of it is a very big (expensive) microscope

∗ Allows one to detect and measure the results of the collisions

Collaboration now consists of

• 39 institutions from US, Japan, Italy, Canada and Taiwan • 450 physicists

Why so many people?

• takes enormous scientific effort to mount experiment

• continual improvements and repairs

In 1995, the Evidence was in...

In about 10 billion collisions, found about 50 events

We observe about the right number of different kinds of events

• Observe 1 μ^{+} μ^{−} event and 5 eμ events

• Expect 1.3 ± 0.3 events from background

• Observe ∼ 40 events with W + ≥ 3 jets

Probability of background looking like this is < 10^{−6}

Whats Happening Now?

Finished taking data in 1996, now completing analysis

• Now have enough top decays to measure its properties – mass (175 ± 6 GeV/c^{2})

– how it decays (like we expect)

– how it gets produced (just about right)

Our group is now improving CDF – ready by 2000

• Collider will start running again at ×10 higher intensity • Collect about more top events

Will be using this data to

• measure properties of top and bottom quarks

• search for new “physics”

Where Do We Go From Here

Still at least one missing piece – Higgs Boson

• We think it is the key for why “mass” exists

• Building a new proton-proton collider in Geneva – Large Hadron Collider

– 7 times more energetic than Fermilab Tevatron – 100 times higher collision rate

– experimental teams have over 1000 scientists

• But don’t hold your breath – Turn-on scheduled for 2005!

This question of Mass is really important!

• Still don’t really understand what energy is

• Gravity remains a mystery

• Don’t really know WHY we have the world that we do

– So why truth and beauty?

∗ Heaviest building blocks in nature

∗ Has to be a clue there (but what?)

So What?

Top quark discovery is a triumph of the Standard Model

• validates 17 years of work

• ends a great deal of speculation

• strong indication that we are on the right track – focus on remaining pieces of puzzle now

We have no idea of the practical benefit this will bring

• basic research is something like Pandora’s Box

• history has shown that basic research eventually has impact – totally unpredictable

– impossible to value accurately

Why Are These Questions So Important?

Humanity seems to be “driven” to wonder

• We seek out frontiers – Physical challenges

– Intellectual challenges

• Is there somethere out there we don’t understand? – science allows us to focus our questions – ask what appear to be the “best” questions

• We question where we fit in?

Push our technological frontier

• This physics is the most “basic” of basic sciences

• Historically, very rich in spin-offs – Nuclear medicine, imaging

– Smoke detectors and other sensors

– Nuclear power, weapons

– Large scale data processing ...... List goes on

Presents some of the greatest intellectural challenges

• Attracts some of our most creative and energetic people • Allows one to leave a “mark” on our world

Page 21 of ?? pages.

Summary

Haven’t reached the end of the story

• Many open questions, for example – What and where is mass?

– What really are quarks? leptons? neutrinos?

– How did our universe become what it is today?

But we are making progress

• We have a “picture” that works pretty well already

• Focusing on those pieces that don’t quite fit

Soon we hope to have a much better understanding of

• the bricks,

• the mortar,

• ...and what nature builds with it.

Consider Now the Quark

Consider now the Quark, which is

A Concept sub-atomical.

No man alive has seen its phiz; Perhaps it isn’t Is at all.

A Concept sub-atomical

Too tenuous I find to prove.

Perhaps it isn’t Is at all.

This goes for Hate, and also Love.

Too tenuous I find to prove

The Sun, the Shadow, and the Wind.

This goes for Hate, and also Love. And other matters of the kind.

The Sun, the Shadow, and the Wind.

The Dream, the Doing, the Depair,

And other matters of the kind I find are proven best in prayer.

The Dream, the Doing, the Despair,

And other matters, being His,

I find are proven best in prayer. Consider now the Quark, which Is.

The Primer of Prosidy - Espy